OpenStudy (anonymous):

A spring bounces up and down according to the model d(t) = 4sin(30t) + 3, where d(t) is the displacement in cm from the rest position and t is time in seconds. What is the equation of the axis? Justify your answer.

OpenStudy (campbell_st):

well the centre of the curve is at d(t) = 3 the amplitude of the curve is 4 the period is \[\frac{2\pi}{30}= \frac{\pi}{15}\] hope it helps